OpenStudy (anonymous):

there are 28 hours left until the final exam on calculus 1...and i havent read anything the entire semester...what can i do to pass?

OpenStudy (amistre64):

take it again next term?

OpenStudy (dumbcow):

28 hrs left?, im guessing this was on online class

OpenStudy (anonymous):

if you have any old papers with answers - cram cram and cram!!

OpenStudy (anonymous):

nope...real life...aprox:28 hrs left tic-toc...

OpenStudy (dumbcow):

how did you do on the prev tests?

OpenStudy (anonymous):

70 and 74...on the midterms...but i didn't study anything EVER!

OpenStudy (anonymous):

CRAM?

OpenStudy (anonymous):

-ster?

OpenStudy (dumbcow):

not too bad considering...practice practice practice and you'll pass http://tutorial.math.lamar.edu/Classes/CalcI/CalcI.aspx this will help

OpenStudy (anonymous):

thanks!...i'll check it out cause the book is a bit confusing...and i cant read my notes since my letters are like hyroglyphs

OpenStudy (anonymous):

cram = constant hard study

OpenStudy (anonymous):

:S ... and where do u derive that from?

OpenStudy (anonymous):

how did u come up with that?

OpenStudy (anonymous):

cram uk slang for a session of hard study in a short time thats all

OpenStudy (anonymous):

let me dig on that...since all i do in my life is wait for the last day to study hard :D

OpenStudy (anonymous):

yea - i used to do that a lot of the time when i was younger!

OpenStudy (anonymous):

first result :) lol ....it tottally describes me!!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how did it work for you?

OpenStudy (anonymous):

hatzitesla is able to study for the calculus exam at the asymptotically increasing rate of \[f(t) = \frac{t}{900-t^2},\] where t is the number of hours studied. If the total proportion of material understood after N hours is given by \[P(N) = \int\limits_{0}^{N} \frac{t}{900-t^2}dt,\] and hatzitesla only needs P(N) >= 100% = 1.0 to ace the exam, what is the minimum number of hours N that hatzitesla needs to spend studying instead of chatting on openstudy?

OpenStudy (anonymous):

holy...!!!man....that was some....math genius pellet....thumbs up!!!...and openstudy is actualy helpfull.(instead of fb for that matter).